# Self-Calibrated CLIP for Training-Free Open-Vocabulary Segmentation

### v2

@@ Abstract
Recent advancements in pre-trained vision-language models like CLIP, have enabled the task of open-vocabulary segmentation. CLIP demonstrates impressive zero-shot capabilities in various downstream tasks that require holistic image understanding. However, due to its image-level pre-training, CLIP struggles to capture local details, resulting in poor performance in segmentation tasks. Our analysis reveals that anomaly tokens emerge during the forward pass, drawing excessive attention from normal patch tokens, thereby diminishing spatial awareness. To address this issue, we propose Self-Calibrated CLIP (SC-CLIP), a training-free method that calibrates CLIP to produce finer representations while preserving its original generalization ability, without introducing new parameters or relying on additional backbones. Specifically, we first identify and resolve the anomaly tokens to mitigate their negative impact. Next, we enhance feature discriminability and attention correlation by leveraging the semantic consistency found in CLIP’s intermediate features. Furthermore, we explore how to effectively employ multi-level feature fusion under the training-free setting. Collectively, these strategies enhance CLIP’s feature representation with greater granularity and coherence. Experimental results demonstrate the effectiveness of SC-CLIP, achieving state-of-the-art results across all datasets and surpassing previous methods by 9.5%. Notably, SC-CLIP boosts the performance of vanilla CLIP ViT-L/14 by 6.8 times. Our source code is available here.
1 Introduction
@@
Open-Vocabulary Segmentation (OVS) is an emerging task in computer vision that aims to segment arbitrary categories based on the textual inputs, overcoming the limitations of predefined category sets. To achieve this, models must generalize beyond the training data. Vision-language pretrained models such as CLIP [41], demonstrate remarkable zero-shot capabilities by leveraging large-scale image-text pairs, effectively fulfilling these requirements. However, CLIP’s image-text matching strategy leads to an excessive emphasis on global context, neglecting local and fine-grained details essential for dense prediction tasks. Consequently, directly applying CLIP to segmentation tasks yields poor performance. For example, as shown in Figure 1, the segmentation result generated by patch-text cosine similarity exhibits considerable noise. CLIP ViT-B/16 achieves only 8.9% mIoU on the COCO-Object dataset [6], significantly lagging behind its ability on image-level recognition.
@@
Figure 1: Left: Vanilla CLIP produces a noisy segmentation map, while our Self-Calibrated CLIP (SC-CLIP) generates a clearer and finer result. Right: Comparison of the open-vocabulary segmentation performance, where SC-CLIP achieves the best results.
@@
To address CLIP’s limitations in capturing local details, recent studies have proposed various modifications to its last layer. One line of research [74, 49, 4, 29, 25, 18, 22, 44] introduces correlative attention, replacing the original Q-K attention with alternatives like K-K, to enhance focus on relevant regions. But these methods still operate on the global and noisy inputs, hindering their effectiveness. Another approach [26, 54] incorporates additional backbones [7, 24, 43] to provide richer spatial details. Despite performance gains, they fail to fully exploit CLIP’s semantic knowledge and impose extra computational costs. In fact, both strategies overlook the underlying causes of CLIP’s global focus, preventing them from fundamentally resolving the issue of diminished spatial awareness in CLIP.
@@
Motivated by this, we begin with an in-depth analysis of CLIP. As indicated by the orange circles in Figure 2 (a), we observe that different patch tokens consistently exhibit high activation regions within their attention map. These regions attract excessive attention from other normal patches, distracting their focus away from local and relevant areas. To further investigate, we apply PCA [1] analysis to CLIP’s features in Figure 2 (b), revealing these over-attended tokens differ significantly from the normal ones (including the [CLS] token). Thus, we refer to them as anomaly tokens. We attribute CLIP’s disadvantage in dense prediction tasks to the emergence of anomaly tokens, which leads to uniform attention activations across locations. This disrupts the attention’s ideal ability to extract relevant semantics, resulting in feature homogenization that diminishes local awareness and further exacerbates noise in the feature maps.
@@
Figure 2: Anomaly tokens in CLIP. (a) visualizes the attention map of various selected patches (marked by $\bigstar$), all exhibit excessive focus on the same regions (indicated by orange circle), aligning with the identified outliers in the PCA analysis in (b).
@@
Building on the analysis, we propose enhancing CLIP’s feature representation by weakening the influence of anomaly tokens. To this end, we introduce Self-Calibrated CLIP (SC-CLIP), a training-free approach that leverages CLIP’s inherent properties for effective calibration, strengthening its perception on local and relevant regions. On the one hand, we propose directly resolving these anomaly tokens. To identify them, we apply the Local Outlier Factor (LOF) algorithm [5], a method for detecting outliers. Once located, we replace these anomaly tokens with values interpolated from their spatial neighbors, considering that spatially close regions often share similar semantics. This not only serves as a regularization to prevent inappropriate attention focus but also reassigns meaningful semantic information to the anomaly tokens, aligning them with the local context. On the other hand, to relieve the feature homogenization problem caused by anomaly tokens, we propose a self-adjusting strategy to enhance feature discriminability and attention correlation. We observe that the intermediate layers’ features of CLIP exhibit good semantic coherence, as shown in Figure 4. In order to retain the rich semantics of deep features and the spatial consistency of intermediate features, we utilize the latter to adaptively aggregate deep features, while simultaneously enhancing attention correlation. This self-adjusting approach improves the overall semantic coherence. Furthermore, we explore how to effectively leverage multi-level feature fusion under the training-free setting and propose a two-pass strategy to enhance the capture of details at different scales.
@@
  • We calibrate CLIP with its inherent properties. Specifically, we propose resolving the anomaly tokens to mitigate their negative impact. And we utilize the spatial consistency in CLIP’s mid-level features to enhance feature discriminability and attention correlation.
@@
Figure 3: Resolving the Anomaly Tokens. (a) Illustration of the resolving process. We plot the feature map using the mean value of each token. After locating the anomaly tokens (the center of red square), we replace them with the interpolated values obtained from their neighboring regions. (b) Effect on the attention map. We highlight the changes for a normal token ($\bigstar$), and an anomaly token ($\blacktriangle$).
2 Related Work
Vision-Language Pre-training.
@@ 2 Related Work
Vision-language models [21, 41, 67, 28, 70, 58] pretrained on large-scale web data, represented by CLIP [41], employ contrastive learning to align images with associated captions, demonstrating remarkable zero-shot capabilities across various downstream tasks that require holistic image understanding [23, 37, 2]. However, CLIP’s image-level pre-training, which relies on the [CLS] token to represent the whole image, causes the model to overly focus on global features at the expense of local and fine-grained details. This limitation hinders its performance in dense prediction tasks which require pixel-level understanding. To address this, our work focuses on effectively adapting CLIP for segmentation task while preserving its original knowledge and cross-modal capabilities.
Open-Vocabulary Segmentation.
@@
The open-vocabulary segmentation (OVS) task focuses on segmenting images with arbitrary text queries by leveraging the zero-shot capabilities of vision-language models [21, 41]. This task has broad potential applications across many vision-related tasks [69, 35, 52, 51, 66, 65, 75, 72]. Existing works can be broadly classified into three categories: fully-supervised, weakly-supervised and training-free. Fully-supervised methods [16, 62, 13, 31, 61, 34, 63, 19, 68, 56, 27, 9, 64, 40] require fine-tuning on pixel-level annotated datasets. Weakly-supervised approaches [39, 42, 8, 59, 36, 60, 71, 20, 39] reduce reliance on dense annotations by using image-text pair to guide region grouping. Training-free methods [74, 49, 4, 29, 26, 54, 22, 44, 18, 3, 53, 33, 47] directly use CLIP for segmentation by making minimal adjustments to the model’s architecture without additional training.
@@
Our approach needs no training and falls into the third category. Recent methods have discovered that CLIP’s final-layer exhibits poor spatial consistency and proposed various modifications. For instance, SCLIP [49] replaces the original Q-K attention with the combination of Q-Q and K-K attention, to enhance correlation. GEM [4] proposes generalized self-self attention and a set of regularizations. ClearCLIP [25] identifies the primary source of noise stems from the residual connections, and proposes to remove it. CLIPTrase [44] proposes using self-self attention along with clustering for post-processing. Other works leverage vision models like DINO [7] and SAM [24] for providing details. For example, CLIP-DINOiser [54] refines feature maps using the affinity learned from DINO’s feature correspondence, while ProxyCLIP [26] applies them to adjust attention weights. In contrast, our approach improves semantic coherence by exploiting CLIP’s internal properties and explicitly resolving the anomaly tokens.
3 Method
@@
In this section, we begin with an overview of CLIP and its dense inference pipeline. We then present SC-CLIP, our approach aimed to enhance CLIP’s representation. First, we identify and mitigate the impact of anomaly tokens. Next, we utilize the mid-level features with spatial consistency to adaptively aggregate the deep features, while simultaneously enhancing attention correlation. Besides, we explore effective multi-level feature fusion to enrich details.
3.1 Preliminaries
@@
CLIP ViT model encodes an input image into a token sequence $\mathbb{X}=[x_{\text{cls}},x_{1},\dots,x_{N}]$, where $x_{\text{cls}}$ is the [CLS] token and the others represent dense visual features, comprising $N$ patch tokens. The CLIP model includes multiple layers, with each layer $l$ processing the input $\mathbb{X}^{(l-1)}$ as follows:
@@
where SA, FFN and LN denote self-attention module, feed-forward network, and layer normalization, respectively.
@@
3.2 Resolving the Anomaly Tokens
We attribute CLIP’s limitations in dense prediction tasks to the presence of anomaly tokens within its features, which deviate significantly from normal tokens. These anomaly tokens cause other tokens to disproportionately focus on them in deep layers, leading to identical attention activation, which undermines attention’s capacity to extract semantically coherent regions. Since attention mechanism [48, 14] manages spatial arrangements, this pattern intensifies noise in the feature map, ultimately degrading performance. Studies [12, 55, 44] suggest that pretrained models like CLIP may identify redundant tokens and use them to gather global information, thereby expediting processing. However, these tokens lack semantics, conveying minimal information about their original positions.
To address the negative impact of anomaly tokens, we propose an intuitive approach to resolve them in $\mathbb{X}^{penul}$ before the last layer, as illustrated in Figure 3 (a). First, it is essential to identify the anomaly tokens. As previously analyzed in Figure 2, they exhibit a clear distinction from other tokens. To detect these anomalies, we employ the Local Outlier Factor (LOF) algorithm [5], a widely used method for anomaly detection. LOF identifies outliers by measuring the local density deviation of a data point relative to its neighbors. Specifically, it computes the local reachability density and assigns a high LOF score to points with significantly lower density than their neighbors, indicating potential anomalies. Besides, we implement the PyTorch-based LOF algorithm to enhance the computational efficiency.
Once the anomalies are located, we replace them with the values interpolated from their $3\times 3$ neighboring regions, based on the assumption that features in spatially adjacent regions are generally similar. Specifically, we apply a $3\times 3$ convolution kernel with the center set to 0. And if neighboring regions contain any other anomaly tokens, they are excluded from the interpolation. The operation’s formula is provided below, where $\tilde{\mathbb{X}}^{penul}$ denotes the feature after operation and $\mathcal{A}$ is the set of anomaly tokens.
$\tilde{\mathbb{X}}_{(x,y)}^{penul}$ $=\frac{\sum_{i=-1}^{1}\sum_{j=-1}^{1}w_{i,j}\cdot\mathbb{X}_{(x+i,y+j)}^{penul}}{\sum_{i=-1}^{1}\sum_{j=-1}w_{i,j}},\quad\forall(x,y)\in\mathcal{A}$
@@ 3.2 Resolving the Anomaly Tokens
We believe that eliminating anomaly tokens offers two key benefits. First, it acts as a form of regularization for the attention. Second, it reassigns semantic information to these anomaly tokens, aligning them with the local context. As shown in the Figure 3 (b), the vanilla attention in CLIP has two issues: 1) normal tokens excessively focus on anomaly tokens, reducing attention on relevant local areas, and 2) anomaly tokens focus on prominent objects and other anomaly tokens. After resolving the anomaly tokens via interpolation, normal tokens refocus on relevant local regions, while the updated anomaly tokens also shift attention to appropriate areas. In summary, our method reduces the impact of anomaly tokens and enhances attention correlation.
3.3 Self-Adjusting for Semantic Coherence

### v3

@@ Abstract
Recent advancements in pre-trained vision-language models like CLIP have enabled the task of open-vocabulary segmentation. CLIP demonstrates impressive zero-shot capabilities in various downstream tasks that require holistic image understanding. However, due to the image-level contrastive learning and fully global feature interaction, ViT-based CLIP struggles to capture local details, resulting in poor performance in segmentation tasks. Our analysis of ViT-based CLIP reveals that anomaly tokens emerge during the forward process, attracting disproportionate attention from normal patch tokens and thereby diminishing spatial awareness. To address this issue, we propose Self-Calibrated CLIP (SC-CLIP), a training-free method that calibrates CLIP to generate finer representations while preserving its original generalization ability—without introducing new parameters or relying on additional backbones. Specifically, we mitigate the negative impact of anomaly tokens from two complementary perspectives. First, we explicitly identify the anomaly tokens and replace them based on local context. Second, we reduce their influence on normal tokens by enhancing feature discriminability and attention correlation, leveraging the inherent semantic consistency within CLIP’s mid-level features. In addition, we introduce a two-pass strategy that effectively integrates multi-level features to enrich local details under the training-free setting. Together, these strategies enhance CLIP’s feature representations with improved granularity and semantic coherence. Experimental results demonstrate the effectiveness of SC-CLIP, achieving state-of-the-art results across all datasets and surpassing previous methods by 9.5%. Notably, SC-CLIP boosts the performance of vanilla CLIP ViT-L/14 by 6.8 times. Furthermore, we discuss our method’s applicability to other vision–language models and tasks for a comprehensive evaluation. Our source code is available at https://github.com/SuleBai/SC-CLIP.
Index Terms:
Open-vocabulary segmentation, Training-free
I Introduction
@@
Open-Vocabulary Segmentation (OVS) is an emerging task in computer vision that aims to segment arbitrary categories based on the textual inputs, overcoming the limitations of predefined category sets. To achieve this, models must generalize beyond the training data. Vision-language pretrained models such as CLIP [52], demonstrate remarkable zero-shot capabilities by leveraging large-scale image-text pairs, effectively fulfilling these requirements. However, the image-level pre-training strategy and fully global feature interactions of ViT-based CLIP lead to an excessive emphasis on global context, neglecting local and fine-grained details essential for dense prediction tasks. Consequently, directly applying ViT-based CLIP to segmentation tasks yields poor performance. For example, as shown in Figure 1, the segmentation result generated by patch-text cosine similarity exhibits considerable noise. CLIP ViT-B/16 achieves only 8.9% mIoU on the COCO-Object dataset [7], significantly lagging behind its ability on image-level recognition.
@@
Figure 1: Left: Vanilla CLIP produces a noisy segmentation map, while our Self-Calibrated CLIP (SC-CLIP) generates a much clearer and finer result. Right: Performance comparison of the open-vocabulary segmentation methods, where our SC-CLIP achieves the best results across all benchmarks.
@@
To address ViT-based CLIP’s limitations in capturing local details, recent studies have proposed various modifications to its last layer. One line of research [86, 62, 5, 39, 33, 23, 30, 55] introduces correlative attention, replacing the original $\mathbf{Q}\mathbf{K}^{\top}$ attention with alternatives like $\mathbf{K}\mathbf{K}^{\top}$, to enhance focus on relevant regions. But these methods still operate on the global and noisy inputs, hindering their effectiveness. Another approach [34, 68] incorporates additional backbones [9, 32, 54] like DINO [9] to provide richer spatial details. Despite performance gains, they fail to fully exploit CLIP’s semantic knowledge and impose extra computational costs. In fact, both strategies overlook the underlying causes of CLIP’s global focus, preventing them from fundamentally resolving the issue of diminished spatial awareness in CLIP.
@@
Motivated by these limitations, we begin with an in-depth analysis of CLIP. As indicated by the orange dashed circles in Figure 2 (a), we observe that different patch tokens consistently exhibit high activation regions within their attention map. These regions attract excessive attention from other normal patches, distracting their focus away from local and relevant areas. To further investigate, we perform PCA [1] on the patch-level features from CLIP’s last layer and project them into a 2D space, as shown in Figure 2 (b), revealing these over-attended tokens significantly differ from the normal ones (including the [CLS] token). Thus, we refer to them as anomaly tokens. We attribute CLIP’s disadvantage in dense prediction tasks to the emergence of anomaly tokens, which leads to uniform attention activations across locations. This disrupts the attention’s ideal ability to extract relevant semantics, resulting in feature homogenization that diminishes local awareness and further exacerbates noise in the feature maps.
@@
Figure 2: Anomaly tokens in CLIP. In (a), we visualize the attention maps of various selected patches (marked by $\bigstar$), which all exhibit excessive focus on the same regions (indicated by the orange dashed circle). And this region aligns with the outliers identified in the PCA analysis shown in (b).
Building on this analysis, we propose enhancing CLIP’s feature representation by weakening the influence of anomaly tokens. To this end, we introduce Self-Calibrated CLIP (SC-CLIP), a training-free approach that leverages CLIP’s inherent properties for effective calibration, strengthening its perception on local and relevant regions.
@@
Specifically, we address the negative impact of anomaly tokens from two perspectives. On the one hand, we propose directly resolving these anomaly tokens. To identify them, we apply the Local Outlier Factor (LOF) algorithm [6], a method for detecting outliers. Once located, we replace these anomaly tokens with values interpolated from their spatial neighbors, considering that spatially close regions often share similar semantics. This not only serves as a regularization to prevent inappropriate attention focus but also reassigns meaningful semantic information to the anomaly tokens, aligning them with the local context. On the other hand, to relieve the feature homogenization problem caused by anomaly tokens, we propose a self-adjusting strategy to enhance feature discriminability and attention correlation. We observe that the mid-layers’ features of CLIP exhibit good semantic coherence, as shown in Figure 4. In order to retain the rich semantics of deep features and the spatial consistency of mid-level features, we utilize the latter to adaptively aggregate deep features, while simultaneously enhancing attention correlation. This self-adjusting approach improves the overall semantic coherence. Furthermore, we explore how to effectively leverage multi-level feature fusion under the training-free setting and propose a two-pass strategy to enhance the capture of details at different scales. The key insights lie in ensuring feature compatibility across layers through alignment with CLIP’s final layer, and preserving the integrity of the last-layer features to maintain strong cross-modal correspondence.
@@
  • We mitigate the negative effects of anomaly tokens from two perspectives. First, we explicitly address the anomaly tokens based on local context. Second, we reduce their impact on normal tokens by enhancing feature discriminability and attention correlation, leveraging the spatial consistency inherent in CLIP’s mid-level features.
@@
Figure 3: Resolving the Anomaly Tokens. (a) Illustration of the resolving process. We plot the feature map using the mean value of each token. After locating the anomaly tokens (the center of red square), we replace them with the interpolated values obtained from their neighboring regions. (b) Effect on the attention map. We highlight the changes on attention map for a normal token ($\bigstar$), and an anomaly token ($\blacktriangle$).
II Related Work
This section focuses on two interrelated areas: vision-language pretrained models and open-vocabulary segmentation. We highlight significant advancements and ongoing challenges, providing a critical overview that identifies gaps in current research and proposes directions for future exploration.
II-A Vision-Language Pretrained Models
@@
Vision-language models [28, 52, 79, 36, 82, 72] pretrained on large-scale web data, represented by CLIP [52], employ contrastive learning to align images with associated captions. These models have demonstrated remarkable zero-shot capabilities across various downstream tasks that require comprehensive image understanding, such as visual question answering and image-text retrieval [31, 47, 2]. However, CLIP’s image-level pre-training, which relies on the [CLS] token to represent the whole image, causes the model to excessively focus on the global features at the expense of local and fine-grained details. This limitation hinders its performance in dense prediction tasks which require pixel-level understanding. To address this, our work focuses on effectively adapting CLIP for segmentation task while preserving its original knowledge and cross-modal alignment capabilities.
II-B Open-Vocabulary Segmentation
@@
The open-vocabulary segmentation (OVS) task [66] focuses on segmenting images with arbitrary text queries by leveraging the zero-shot capabilities of vision-language models [28, 52]. This task has broad potential applications across many vision-related tasks [83, 78, 65, 26, 81, 45, 59, 37, 84]. Existing works can be broadly classified into three categories: fully-supervised, weakly-supervised and training-free. Fully-supervised methods [63, 21, 76, 18, 41, 44, 75, 24, 80, 70, 35, 14, 77, 51, 87, 38] require fine-tuning on pixel-level annotated datasets. Weakly-supervised approaches [49, 53, 10, 73, 46, 74, 25, 12, 49] reduce reliance on dense annotations by using image-text pair to guide region grouping. Training-free methods [86, 62, 5, 39, 34, 68, 30, 55, 23, 3, 67, 43, 58] directly use CLIP for segmentation by making minimal adjustments to the model’s architecture without additional training.
@@
Our approach needs no training and falls into the third category. Recent methods have discovered that CLIP’s final layer exhibits poor spatial consistency and proposed various modifications. For instance, SCLIP [62] replaces the original $\mathbf{Q}\mathbf{K}^{\top}$ attention with the combination of $\mathbf{Q}\mathbf{Q}^{\top}$ and $\mathbf{K}\mathbf{K}^{\top}$ attention, to enhance correlation. GEM [5] proposes generalized self-self attention and a set of regularizations. ClearCLIP [33] identifies that the primary source of noise stems from the residual connections and proposes to remove it. CLIPTrase [55] notices the [CLS] token may disrupt the patch correlations and proposes using self-self attention along with clustering and denoising for post-processing. Other works leverage additional vision models like DINO [9] and SAM [32] for providing fine-grained spatial details. For example, CLIP-DINOiser [68] refines feature maps using the affinity learned from DINO’s feature correspondence, while ProxyCLIP [34] applies them to adjust attention weights. In contrast, our approach improves semantic coherence by exploiting CLIP’s internal properties and explicitly resolving the negative impact of anomaly tokens.
III Method
@@
In this section, we begin with an overview of CLIP and its dense inference pipeline. We then present SC-CLIP, our training-free approach designed to enhance CLIP’s dense representation. We address anomaly tokens from two perspectives. First, we directly identify these tokens and replace them based on local context. Next, to mitigate their influence on normal tokens, we leverage mid-level features with stronger spatial consistency to guide the adaptive aggregation of deep features and enhance the attention correlation. Furthermore, we propose a two-pass strategy to effectively integrate multi-level features and enrich spatial details.
III-A Preliminaries
@@
The CLIP ViT model encodes an input image into a token sequence $\mathbb{X}=[x_{\text{cls}},x_{1},\dots,x_{N}]$, where $x_{\text{cls}}$ is the [CLS] token and the others represent dense visual features, comprising $N$ patch tokens. The CLIP model includes multiple layers, with each layer $l$ processing the input $\mathbb{X}^{(l-1)}$ as follows:
@@
where SA, FFN, and LN denote the self-attention module, feed-forward network, and layer normalization, respectively.
@@
III-B Resolving the Anomaly Tokens
We attribute CLIP’s limitations in dense prediction tasks to the presence of anomaly tokens within its features, which deviate significantly from normal tokens. These anomaly tokens cause other tokens to disproportionately focus on them in deep layers, leading to identical attention activation, which undermines attention’s capacity to extract semantically coherent regions. Since attention mechanism [61, 19] manages spatial arrangements, this pattern intensifies noise in the feature map, ultimately degrading performance. Studies [17, 69] suggest that pretrained models like CLIP may identify redundant tokens and use them to gather global information, thereby expediting processing. However, these tokens lack semantics, conveying minimal information about their original positions. Existing methods do not explicitly address the anomaly tokens or the impact they cause to other normal tokens.
To address the negative impact of anomaly tokens, we propose an intuitive approach to directly resolve them in $\mathbb{X}^{penul}$ before the last layer, as illustrated in Figure 3 (a). First, it is essential to identify the anomaly tokens. As previously analyzed in Figure 2, they exhibit a clear distinction from other tokens. To detect these anomalies, we employ the Local Outlier Factor (LOF) algorithm [6], a widely used method for anomaly detection. LOF identifies outliers by measuring the local density deviation of a data point relative to its neighbors. Specifically, it computes the local reachability density and assigns a high LOF score to points with significantly lower density than their neighbors, indicating potential anomalies. Besides, we implement the PyTorch-based LOF algorithm to enhance the computational efficiency.
Once the anomalies are located, we replace them with the values interpolated from their $3\times 3$ neighboring regions, based on the assumption that features in spatially adjacent regions are generally similar. Specifically, we apply a $3\times 3$ convolution kernel with the center set to 0. And if neighboring regions contain any other anomaly tokens, they are explicitly excluded from the interpolation. The operation’s formula is provided below, where $\tilde{\mathbb{X}}^{penul}$ denotes the feature after operation and $\mathcal{A}$ is the set of anomaly tokens.
$\tilde{\mathbb{X}}_{(x,y)}^{penul}$ $=\frac{\sum_{i=-1}^{1}\sum_{j=-1}^{1}w_{i,j}\cdot\mathbb{X}_{(x+i,y+j)}^{penul}}{\sum_{i=-1}^{1}\sum_{j=-1}^{1}w_{i,j}},\quad\forall(x,y)\in\mathcal{A}$
@@ III-B Resolving the Anomaly Tokens
We believe that eliminating anomaly tokens offers two key benefits. First, it acts as a form of regularization for the attention. Second, it reassigns semantic information to these anomaly tokens, aligning them with the local context. As shown in the Figure 3 (b), the vanilla attention in CLIP has two issues: 1) normal tokens excessively focus on anomaly tokens, reducing attention on relevant local areas, and 2) anomaly tokens focus on prominent objects and other anomaly tokens. After resolving the anomaly tokens via interpolation, normal tokens refocus on relevant local regions, while the updated anomaly tokens also shift attention to appropriate areas.
III-C Self-Adjusting for Semantic Coherence